Morning Litho Status Report 10/25

Gary Sosa gsosa at stanford.edu
Mon Oct 25 07:30:44 PDT 2010


Good Morning All.  Here is the Litho Status Report.

  Karl Suss#2- User reports a theta rotation problem. Will investigate 
and repair as needed
  SVG Coat #1-  User reports no EBR dispensing. Will check out system 
and repair as needed
  SVG Coat #2- User reports that wafers not transferring from sender to 
VP. Will check out and repair as needed.

   Please check back in Coral for the current status of tools.


    Thanks.... The Litho Team


*/Photolithography Area Status/*

	

	

*/Monday, October 25, 2010/*

	

	

	

	

*/General Information / Lab Activities/*

	

	

	

EVG Bonder Back to Quartz pressure plate today

*/Exposure Tools/**//*

	

Status

	

Comments

	

Priority

ASML Stepper

	

*UP*

	

	

EVAlign

	

*UP*

	

	

EVAlign2

	

*UP*

	

	

EVG-Imprint

	

*UP*

	

	

Karl Suss 1

	

*UP*

	

	

Karl Suss 2

	

*PROBLEM*

	

Theta rotation problem

	

3

*/Manual Coaters/*

	

Status

	

	

EVGSpraycoat

	

*PROBLEM*

	

Wobbly chuck

	

4

Headway

	

*UP*

	

	

Laurel

	

*UP*

	

	

*/Coaters/*

	

Process Status

	

	

SVGCoat

	

3612 Process

	

*Problem*

	

No EBR dispensing

	

1

3617 Process

	

*Problem*

	

	

220-7 Process

	

*Problem*

	

Optimizing process

	

SVGCoat2

	

3612 Process

	

*Problem*

	

Transfer problem from sender to VP

	

2

955-0.7 Process

	

*Problem*

	

	

220-3 Process

	

*Problem*

	

	

*/Developers/*

	

Status

	

	

SVGDEV

	

*UP*

	

	

SVGDEV2

	

*UP*

	

	

*/Ovens / Hotplates/*

	

Status

	

	

BLUEM

	

*UP*

	

	

Blue M 90 C.

	

*UP*

	

	

Blue M 110 C

	

*PROBLEM*

	

Temperature Drift. Controller on order

	

6

Blue M 150 C

	

*PROBLEM*

	

Temperature Drift. Controller on order

	

7

Fusion

	

*DOWN*

	

Hotplate experiments

	

UV-Cure

	

*UP*

	

	

White-Oven

	

*UP*

	

	

YES Vapor Prime

	

*UP*

	

	

*/Hotplates- /*

	

	

	

Laurel-left

	

*UP*

	

	

Laurel- right

	

*UP*

	

	

Headway

	

*UP*

	

	

Torrey Pines- left

	

*UP*

	

	

90 Deg C

	

*UP*

	

	

115 Deg C

	

*UP*

	

	

Torrey Pines- Right

	

*UP*

	

	

*/Metrology/Inspection/*

	

Status

	

	

Alpha-Step 500

	

*UP*

	

	

Nanospec 2-Litho

	

*UP*

	

	

Nanospec 1-Etch

	

*UP*

	

	

P2 Profiler

	

*PROBLEM*

	

Error -11 on Unload, reset to recover

	

5

*/Microscpoes-/*

	

	

	

#1 Leica/Analytical

	

*UP*

	

	

#2 Ergolux/Drytek2

	

*UP*

	

	

#3 Olympus/Litho

	

*PROBLEM*

	

Camera Apps software issue

	

8

#4 Nikon Optiphot/Litho

	

*UP*

	

	

#5 Nikon/Litho

	

*UP*

	

	

#6 Leitz/Bosch

	

*UP*

	

	

#7 Leitz/Diffusion

	

*UP*

	

	

#8 Olympus/Litho

	

*UP*

	

	

#9 Olympus/Litho

	

*UP*

	

	

*/Miscellaneous/*

	

Status

	

	

EVBond

	

*PROBLEM*

	

1 bond chuck out for rebuild by EVG

	

KSBonder

	

*UP*

	

	

KSCoat

	

*UP*

	

	

Litho SRD

	

*UP*

	

	

	

	

	

	

	

	


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/litho/attachments/20101025/5c93cb2d/attachment.html>


More information about the litho mailing list